OpenStudy (anonymous):

solve using systems of equations At a football game, tickets cost #3 for the students and $7 for adults. if a total of 1,820 people attended the game and $8000 was collected by ticket sales how many of each type of ticket was sold

OpenStudy (anonymous):

3s+7a=8000 s+a=1820 where s is number of students, a is number of adults. Let me know if you need help solving

OpenStudy (anonymous):

yes i do sorry i get all lost

OpenStudy (anonymous):

k do you want me to explain how i got the equations or does that make sense

OpenStudy (anonymous):

it makes sense when i see it

OpenStudy (anonymous):

k. then to solve we multiply the bottom equation by -3 3s+7a=8000 -3s-3a=-5460 now add the two 4a=2540 a=635 plug back in to get s s+635=1820 s=1185

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

yes that is great thank you a bunch

OpenStudy (anonymous):

no problem